OpenStudy (rebeccapink6398):

@ganeshie8 plz help :D How much is a 15% tip for a meal costing $32? A. $4.80 B. $4.70 C. $2.13 D. $1.50 I think its A. But im not sure... :)

OpenStudy (anonymous):

A

OpenStudy (rebeccapink6398):

I need help not answer.. that's not the point of this website...

OpenStudy (anonymous):

It's A, 4.80 10% of 32$ is $3.20 1/2 of that (ie another 5% is 1.60 3.20 + 1.60 = 4.80

OpenStudy (anonymous):

\[\huge \frac{15}{100} \times 32= 0.15 \times 32 = 4.8\]

OpenStudy (rebeccapink6398):

THX!! :D

OpenStudy (anonymous):

Hence otion A is the correct answer...